# Semantic Mapping in Indoor Embodied AI – A Comprehensive Survey and Future Directions

### v1

@@ Abstract
Intelligent embodied agents (e.g. robots) need to perform complex semantic tasks in unfamiliar environments. Among many skills that the agents need to possess, building and maintaining a semantic map of the environment is most crucial in long-horizon tasks. A semantic map captures information about the environment in a structured way, allowing the agent to reference it for advanced reasoning throughout the task. While existing surveys in embodied AI focus on general advancements or specific tasks like navigation and manipulation, this paper provides a comprehensive review of semantic map-building approaches in embodied AI, specifically for indoor navigation. We categorize these approaches based on their structural representation (spatial grids, topological graphs, dense point-clouds or hybrid maps) and the type of information they encode (implicit features or explicit environmental data). We also explore the strengths and limitations of the map building techniques, highlight current challenges, and propose future research directions. We identify that the field is moving towards developing open-vocabulary, queryable, task-agnostic map representations, while high memory demands and computational inefficiency still remaining to be open challenges. This survey aims to guide current and future researchers in advancing semantic mapping techniques for embodied AI systems.
@@ 1 Introduction
Over the past few years there has been a growing interest across computer vision, natural language and the robotics community in embodied AI – where we study how intelligent agents with an embodiment (e.g. robots, autonomous vehicles) learn to perform tasks through interaction with its environment (Deitke et al., 2022; Puig et al., 2023; Batra et al., 2020a). This paves the way towards creating service robots (Hawes et al., 2017; Khandelwal et al., 2017; Veloso et al., 2015) that can be safely deployed in familiar environments, co-existing with humans and performing various tasks autonomously. The key difference between embodied AI and robotics is that the former focuses on building robot intelligence by interacting with simulated physical world while abstracting most of the low-level control including noisy sensors and actuators. This allows the embodied AI researchers to focus on complex semantic challenges such as object search in unfamiliar environments based on natural language instructions, spatial reasoning, multi-agent systems, interactive object search and much more. Similar to robotics, this requires the agent to possess a blend of sensorimotor skills, understanding of the environment and decision-making abilities, thus enabling them to navigate, manipulate objects and perform complex tasks in the world. Imagine a robot is tasked with finding an object in an unseen environment – “a red and blue striped zebra toy in the nursery”. This is a complex task which requires the robot to have multiple skills – visual perception to identify the toy (Matthias Minderer, 2023), natural language understanding to make sense of the given language instruction (Anderson et al., 2018b), navigation to move to the nursery (Yadav et al., 2022), maintaining a semantic map to remember where the zebra toy was if it has seen it already (Raychaudhuri et al., 2023), reasoning and planning to take actions in order to complete the task (Gordon et al., 2018), and so on.
Among these skills, progressively building and maintaining a map of the world has been found to be especially crucial when the robot is in unfamiliar environments such as in autonomous driving (Bao et al., 2023), search-and-rescue operations (Gautham et al., 2023), automated vacuum cleaning robots (Singh et al., 2023) and others. An accurate map of the environment allows the agent to make informed decisions, handle unexpected situations and perform a complex task autonomously without human supervision. The significance of mapping in robotics navigation methods is motivated by findings in cognitive science (Tolman, 1948; Trullier & Meyer, 2000) which shows that humans and animals create an internal representation of their surroundings in the form of ‘cognitive maps’ or ‘mental maps’ to aid spatial cognition and form navigation strategies. Moreover such cognitive maps allow them to remember and recall the location of objects and places in the environment. Following this, most robotics navigation models are map-based (Kostavelis & Gasteratos, 2015; Filliat & Meyer, 2003; Meyer & Filliat, 2003; Song et al., 2024) and can be broken down into three processes – (a) mapping or memorizing appropriate representation of the environment, (b) localization or determining the current position of the agent on the map, and (c) path planning or choosing a set of actions that lead to the goal, given the map and the current location. While path planning is dependent on the first two, mapping and localization are also dependent on each other. In other words, estimating the current agent position depends on the map and a map can be built once the agent knows where it is located. This problem has been studied traditionally in robotics as simultaneous localization and mapping (SLAM), which is the problem of mapping an unknown environment and at the same time estimating a robot’s pose within it. The robotics community has seen tremendous progress in SLAM approaches which do not have to rely on external reference systems like GPS and instead use the onboard sensors to map a complex environment. This is particularly useful in indoor spaces where GPS is not available. Localization and path planning are beyond the scope of this survey. We instead focus on building maps progressively and memorizing features of an unknown environment while the agent navigates.
Figure 1: Semantic mapping. To perform a complex task in an indoor environment, the robotic agent must possess multiple skills of language understanding, visual perception, navigation, etc. Among these the most crucial is building and maintaining a semantic map of the environment so that it can come back to it while performing the task.
In addition to studying how to build intelligent agents, research in robotics has to consider various low-level aspects (low-level path planning and control, hardware sensors, robot hardware, etc.). In contrast, embodied AI research can focus more on high-level task planning by abstracting out the low-level details. This has led embodied AI researchers to explore map building techniques as part of the high-level task planning and address questions such as ‘is mapping even necessary’ (Partsey et al., 2022), ‘what should be the structure of the map’, ‘what type of information to store inside the map’, and ‘which type of information is useful for what tasks’. There has also been a recent shift in focus to build general-purpose AI solutions by leveraging foundation models (Radford et al., 2021; Oquab et al., 2023; OpenAI, 2023) which has allowed the community to explore building general-purpose, open-vocabulary semantic maps independent of the downstream tasks. These open-vocabulary maps can be later queried using natural language (Gu et al., 2023; Peng et al., 2023; Chen et al., 2023a) or images.
Unlike existing surveys in embodied AI, which often focus on general task advancements (Pfeifer & Iida, 2004; Duan et al., 2022; Deitke et al., 2022), or specific sub-fields like visual navigation (Zhu et al., 2021; Zhang et al., 2022; Wu et al., 2024; Lin et al., 2024) or manipulation (Batra et al., 2020a; Zheng et al., 2024), this paper offers the first comprehensive review of semantic map-building techniques tailored to indoor environments in embodied AI. We organize existing approaches along two key axes – how the maps are structured and what types of information or encoding is stored inside them. By analyzing the advantages and limitations of various methods, we identify gaps in the current literature and propose future research directions to guide the community. This survey aims to unify disparate approaches to semantic mapping in embodied AI, shedding light on its foundational role in enabling intelligent behavior. Beyond merely summarizing existing work, we hope to inspire new research that pushes the boundaries of semantic map building in embodied agents. To this end, we begin with a background on embodied AI and SLAM in Sec. 2, before delving into key questions and challenges of semantic mapping in Sec. 3. We then explore map structures (Sec. 4) and encoding techniques (Sec. 5). Finally, we discuss applications and evaluation techniques in Sec. 6 and conclude with insights into future directions for this rapidly evolving field (Sec. 7). Note that semantic mapping in robotics (real world robots) is out of scope for this survey and we direct the readers to Thrun (2003); Kostavelis & Gasteratos (2015); Lluvia et al. (2021); Racinskis et al. (2023); Sousa et al. (2023) for a comprehensive reading. That said, in this survey we provide comparisons to similar mapping techniques from robotics, wherever applicable.
2 Background Reading
In this section we provide a brief overview of different types of embodied AI tasks (Sec. 2.1) and various approaches towards solving them, including end-to-end (Sec. 2.2) and modular (Sec. 2.3) approaches. As a full survey on Embodied AI tasks is beyond the scope of this survey, please see (Deitke et al., 2022) for a more detailed survey on the tasks and their current state of research. In Sec. 2.4, we discuss classical SLAM based techniques. This discussion aims to provide a background to the readers about the general advance in the embodied AI research and its connection to traditional SLAM-based techniques before we dive in to discuss semantic mapping.
2.1 Embodied AI tasks
Embodied AI tasks vary depending on the type interaction of an agent with its environment. Broadly, we can group embodied tasks into three groups – Exploration task (Chaplot et al., 2019) requires an agent to efficiently explore its environments; Navigation task (Wijmans et al., 2019; Batra et al., 2020b) requires the agent to take actions in order to move around the environment; Manipulation task (Szot et al., 2021; Weihs et al., 2021) requires the agent to perform interactive actions to change the state of other objects in the environment.
The taxonomy of tasks can be further broken down by the target specification provided to the agent, which impacts the information that need to be retained. For instance, for navigation, the following are commonly studied. In Point-Goal Navigation (Wijmans et al., 2019) (PointNav), the agent is given a target coordinate relative to its starting position, whereas in Image-Goal Navigation (ImageNav) it is given a target image (Chaplot et al., 2020b). In the Object-Goal Navigation (ObjectNav) task, the agent needs to navigate to any instance of an object category (Yadav et al., 2022). An extension to the ObjectNav task is the Multi-Object Navigation (MultiON) (Wani et al., 2020) task where the agent is required to navigate to multiple objects in a particular sequence. Vision-and-Language Navigation (VLN) (Anderson et al., 2018b) requires the agent to find the target as specified by a natural language instruction. In Audio-Visual Navigation task, the agent needs to navigate to an object emitting a particular sound in an indoor environment (Chen et al., 2020a; Gan et al., 2019). Depending on the type of task, it may be sufficient to store just the object category (e.g. for ObjectNav) in the map, or it is may be necessarily to retain more finer-grained information (e.g. VLN). In this survey, we will mainly focus on recent work on room-scale map building for navigation as these methods can be extended for maps for manipulation and used for exploration.
2.2 End-to-end approaches
The embodied AI community has seen a lot of progress in training task-specific end-to-end models with reinforcement learning (RL) that directly learns to predict discrete (Wani et al., 2020) or continuous actions (Kalapos et al., 2020) from visual observations (see Fig. 2). These methods may consist of an unstructured memory such as LSTM Dobrevski & Skočaj (2021). However, such representation lacks reasoning about 3D space and geometry and fail to perform well in long-horizon path planning. This has lead to the development of approaches that build an intermediate map representation. Such a map can be implemented using differentiable operations so as to facilitate end-to-end training. Gupta et al. (2017) shows that an egocentric map built this way is beneficial for both PointNav and ObjectNav tasks, whereas Henriques & Vedaldi (2018) learns a global map for the task of localization. While these methods are trained using supervised learning, Wani et al. (2020) use RL to learn to predict actions based on an intermediate global map of the environment to address the complex MultiON task. However, irrespective of the map representation or the mode of training, these approaches need to be retrained every time the task definition changes and even the basic skills required to perform a task need to be learned from scratch.
2.3 Modular approaches
Methods Task End-to-End Modular
Visual Encoder Map Exploration Planner
SemExp (Chaplot et al., 2020a) ObjectNav MaskR-CNN (He et al., 2017) occupancy + explored + semantic labels learned policy Fast Marching
ModLearn (Gervet et al., 2023) ObjectNav MaskR-CNN (He et al., 2017) occupancy + explored + semantic labels learned SemExp Fast Marching
MOPA (Raychaudhuri et al., 2023) MultiON FasterRCNN (Ren et al., 2015) semantic labels Uniform Sampling Exploration PointNav(Wijmans et al., 2019)
CMP (Gupta et al., 2017) PointNav, ObjectNav
MapNet (Henriques & Vedaldi, 2018) Localization
MultiON (Wani et al., 2020) MultiON
Another line of work explores how to breakdown a complex task into a set of basic skills that the agent needs to acquire. Such skills can then be learned independently of each other so that they can be leveraged across various tasks without the need to be retrained from scratch. This has led to various works on modular pipelines (Chaplot et al., 2019; 2020a; Gervet et al., 2023; Raychaudhuri et al., 2023), where each module is responsible for a particular skill and the modules interact with each other to perform the entire task. In the modular approach, it is common to have a visual encoder that processes and encodes the visual information at each time step, a mapper that aggregates the encoded information into a map, a exploration module that determine what parts of the environment needs to be explored, and a planner that determines the low-level action to take. Fig. 2 shows the difference between the end-to-end approach and the module approach, and we elaborate on the modules and popular design choices below.
Visual Encoder.
This module encodes agent observations to produce semantic visual features and predictions at every time step. Prior works have used visual features from pretrained backbones such as ResNet He et al. (2016) or ViT (Dosovitskiy et al., 2020), and often leveraging object detectors MaskRCNN He et al. (2017) or FasterRCNN Ren et al. (2015). As we will see in Sec. 5.2.2, with the development of large pretrained vision-language models and open-vocabulary detectors, pretrained models such as CLIP (Radford et al., 2021), LSeg (Li et al., 2022), DINO (Caron et al., 2021; Oquab et al., 2023), and others are increasingly popular as the basis for building large to open-vocabulary maps. The visual encoder used will determine the information captured in the features, and whether there are detected object instance bounding boxes or segmentations for integration into the mapping modules.
Exploration. This module enables the agent to explore its environment efficiently to either ensure the map is complete (by maximizing the covered area) or selecting unvisited areas where the target is likely to be. Typically, the exploration module selects a point or region to explore given the obstacle map built by the mapper and the current agent pose. Agents can use simple heuristics-based methods such as sampling a point at uniform (Zhang et al., 2021; Raychaudhuri et al., 2023), systematically sampling four corners of a grid centered at the agent (Luo et al., 2022) or selecting a point from the unexplored frontier (Yamauchi, 1997). To decide which frontier point the agent should explore, various strategies are employed, such as selecting the nearest point to the agent (Gervet et al., 2023) or the most promising point based on semantic reasoning. In the semantic reasoning based exploration methods, the agents may select the highest text-image relevance score (Gadre et al., 2023; Yokoyama et al., 2023) from a pretrained large vision-language model such as BLIP-2 (Li et al., 2023), select the highest probabilistic output of the VLM directly (Ren et al., 2024), or leverage a LLM to extract common-sense knowledge (Zhou et al., 2023). Researchers have also used learned policies (Chaplot et al., 2019; 2020a), where the agents are generally trained with RL using rewards, such as coverage (Chen et al., 2019) or curiosity (Pathak et al., 2017; Mazzaglia et al., 2022). Although there is less hand-crafted rules in learning-based methods, they need millions of training steps and careful reward engineering.
Planner. Once a map is built, a low level path-planning module is used to plan a path to the goal location from the agent’s current location. The path consists of low-level actions that can be executed by the agent to move to the goal. While this is implemented as a heuristics-based Fast Marching Method (Sethian, 1996) in most of the prior works, a recent approach MOPA by (Raychaudhuri et al., 2023) has used a learned PointNav policy trained offline with DD-PPO (Wijmans et al., 2019).
2.4 Active SLAM
Figure 3: Active SLAM. At the core of classical mobile robotics lie three core tasks – mapping, localization, and planning. These are often interdependent on each other and overlap to form other tasks such as SLAM (localization and mapping), exploration (mapping and planning), active localization (planning and localization) and active SLAM (mapping, localization, and planning). Figure reproduced from Fairfield (2009).
In this section, we first discuss a popular approach in classical robotics known as active SLAM, with a focus on mapping later. The key idea behind Active SLAM is that, instead of just passively collecting data and using it to construct a map, the robot makes decisions about where to move next in order to reduce uncertainty in the map and its estimated location. This is crucial for an autonomous robot to efficiently perform complex tasks in an unknown environment and involves three key problems (Makarenko et al., 2002; Fairfield, 2009) – mapping, localization, and planning (Fig. 3). In the context of classical robotics, we first define these three problems. Mapping involves collecting sensor data to create an environmental model, accounting for sensor noise and uncertainty. In traditional robotics, these methods address such uncertainties, while embodied AI simulators often assume ideal, noiseless sensors. Localization is the process of determining the robot’s position within the map using sensor data, which can be noisy. Planning is deciding the robot’s next actions, considering constraints like safety and uncertainties in the map, localization, and action outcomes. Note that path planning is a sub-task of planning, focusing on finding the safest and most efficient route to a goal.
These problems are interconnected, leading to tasks such as Simultaneous Localization and Mapping (SLAM), where mapping and localization depend on each other, classical exploration where the robot takes actions to maximize environment coverage and build a complete map, active localization where the robot actively plans and refines its position by taking actions. At the intersection of all three is Active SLAM, where the robot actively reduces uncertainty by exploring, building a map, and localizing itself while refining both. This enables the robot to perform tasks more efficiently.
Next we look at the maps built during active SLAM. Traditionally, the map captured purely geometric information about an environment, by using measurements from LIDAR or depth sensors to build an obstacle map while moving around in an unseen environment. However, this fails to capture crucial semantic cues about the environment needed to perform complex tasks. This has led to the development of Semantic SLAM approaches which additionally detect and identify objects in the scene. Classical methods in Semantic SLAM rely on feature extraction (SIFT (Lowe, 2004), SURF (Bay et al., 2008), ORB (Rublee et al., 2011)) which is then referenced against a dictionary based on Bag-of-Visual Words (Peng & Li, 2016) to determine its closest similarity match. Chen et al. (2022) and Placed et al. (2023) survey the landscape of classical and semantic SLAM approaches in a greater detail. Modern semantic SLAM systems, however, use deep learning based approaches to capture semantic information from its environment. This survey focuses on the mapping approaches in embodied AI with noiseless sensors and perfect odometry, thus allowing us to focus on the built map representation and encoding without having to worry about dealing with noise.
Figure 4: Structure and encoding. The map corresponding to a physical environment can be structured as a topological map (with nodes and edges), a spatial grid, a point-cloud or a hybrid map combining two or more of the others. These structures can store either explicit or implicit information corresponding to the observation made at that location. The figure shows examples of types of explicit information that can be stored at each cell or node. Implicit features are typically extracted and aggregated using vision encoders such as ResNet for close-set vocabulary, or large-vision language models such as CLIP Radford et al. (2021) for open-vocabulary semantics.
@@ 2.4 Active SLAM
Encoding Grid Topological Point Cloud
@@
Implicit (V) CMP (Gupta et al., 2017) MapNet (Henriques & Vedaldi, 2018) MultiON (Wani et al., 2020) SPTM (Savinov et al., 2018) NTS (Chaplot et al., 2020b) CMTP (Chen et al., 2021) VGM (Kwon et al., 2021) $\boxed{t}$
Implicit (VL) CoW (Gadre et al., 2023) VLMap (Huang et al., 2023a) NLMap (Chen et al., 2023a) RoboHop (Garg et al., 2024) OpenScene (Peng et al., 2023) CLIPFields (Shafiullah et al., 2023) CLIP2Scene (Chen et al., 2023c) 3D aware ObjNav (Zhang et al., 2023a)
ConceptGraphs (Gu et al., 2023)

### v2

# Semantic Mapping in Indoor Embodied AI – A Survey on Advances, Challenges, and Future Directions

@@ Abstract
Intelligent embodied agents (e.g. robots) need to perform complex semantic tasks in unfamiliar environments. Among many skills that the agents need to possess, building and maintaining a semantic map of the environment is most crucial in long-horizon tasks. A semantic map captures information about the environment in a structured way, allowing the agent to reference it for advanced reasoning throughout the task. While existing surveys in embodied AI focus on general advancements or specific tasks like navigation and manipulation, this paper provides a comprehensive review of semantic map-building approaches in embodied AI, specifically for indoor navigation. We categorize these approaches based on their structural representation (spatial grids, topological graphs, dense geometric or hybrid maps) and the type of information they encode (implicit features or explicit environmental data). We also explore the strengths and limitations of the map building techniques, highlight current challenges, and propose future research directions. We identify that the field is moving towards developing open-vocabulary, queryable, task-agnostic map representations, while high memory demands and computational inefficiency still remaining to be open challenges. This survey aims to guide current and future researchers in advancing semantic mapping techniques for embodied AI systems.
@@ 1 Introduction
Figure 1: Semantic maps. The survey categorizes semantic map building methods in embodied agents based on their structure and the encoding it stores. Structure: a map of a physical environment can be structured as a topological map (with nodes and edges), a spatial grid, a dense geometric map or a hybrid map combining two or more of the others. Encoding: the structured maps can store either explicit (occupancy, object type, etc.) or implicit information (learned neural features) corresponding to the observation made at that location.
Intelligent agents, whether physical robots or virtual embodied systems, must operate in complex, unstructured environments where effective behavior demands more than just low-level sensing and actuation. To act meaningfully, agents must form structured internal representations that link perception to reasoning and decision-making. Semantic maps serve this role by encoding not only spatial geometry but also high-level semantics of the environment (object categories, affordances, etc.). Semantic mapping is thus foundational in both robotics and embodied AI, particularly in open-world settings like autonomous driving (Bao et al., 2023; Li et al., 2024b), search-and-rescue operations (Gautham et al., 2023), automated cleaning robots (Singh et al., 2023) among others. While traditional mapping techniques prioritize geometric accuracy for localization and obstacle avoidance, recent progress in deep learning, computer vision and multi-modal perception has shifted focus toward semantically rich maps. Despite extensive literature, semantic mapping remains an open and rapidly evolving research area.
Existing surveys review semantic mapping within the context of its usage in downstream applications, thus largely centered on task advancement (Pfeifer & Iida, 2004; Kostavelis & Gasteratos, 2015; Cadena et al., 2017; Xia et al., 2020; Duan et al., 2022; Deitke et al., 2022; Zhu et al., 2021; Zhang et al., 2022b; Wu et al., 2024; Lin et al., 2024; Garg et al., 2020; Achour et al., 2022; Zheng et al., 2024a). Song et al. (2025) has recently reviewed semantic mapping in indoor robotics applications, but focused on the type of semantics acquired and the techniques of such extraction. In contrast, our survey offers a comprehensive review of semantic map-building methods, organized around the underlying map representations themselves, independent of specific downstream tasks. This perspective is both timely and necessary, given the recent advances in foundation models and the growing demand for general-purpose, open-vocabulary, task-agnostic representations.
To provide a principled understanding of semantic map-building methods, this survey categorizes the literature along two fundamental axes (see Fig. 1): map structure (e.g., topological graphs, spatial grids, dense geometric, and hybrid representations) and semantic encoding (explicit annotations vs. learned implicit features). This taxonomy reflects core design choices that influence map scalability, interpretability, generalization, multi-modal fusion and querying. By organizing approaches in this manner, we aim to unify diverse strands of research (see Tab. 1 for a summary of the reviewed papers), highlight the trade-offs between different representations and present key challenges and future opportunities in semantic mapping. This survey centers on semantic mapping approaches in the context of indoor mobile robots, a domain that offers a well-defined, practically relevant and technically rich environment for research. We focus on semantic mapping within embodied AI, that enables controlled environments capable of isolating high-level reasoning abilities of semantic maps from the complexities of noisy sensors and real-world observations. At the same time the survey establishes connections to SLAM-based approaches to underscore their methodological overlap and to bridge foundational techniques in robotics with emerging paradigms in embodied AI.
This survey is organized as follows. We begin the survey in Sec. 2 with background context to ground our discussion on semantic mapping. Sec. 3 introduces semantic maps and lays out the foundational components. We then explore different map structures in Sec. 4, followed by encoding strategies in Sec. 5. Sec. 6 reviews existing evaluation methodologies. In Sec. 7, we examine the key challenges in building semantic maps, leading into Sec. 8, which outlines promising directions for future research. We conclude in Sec. 9 by synthesizing the main insights and takeaways from the survey.
@@
Encoding Grid Topological Dense geometric map
@@
[…]
Implicit (V) CMP (Gupta et al., 2017) MapNet (Henriques & Vedaldi, 2018) MultiON (Wani et al., 2020) RNR-Map (Kwon et al., 2023) SPTM (Savinov et al., 2018) NTS (Chaplot et al., 2020b) CMTP (Chen et al., 2021) VGM (Kwon et al., 2021) $\boxed{t}$ KPConv (Thomas et al., 2019) PointResNet (Desai et al., 2022) Neural Fusion (Mazur et al., 2023)
Implicit (VL) CoW (Gadre et al., 2023) VLMap (Huang et al., 2023a) Le-RNR-Map (Taioli et al., 2023) NLMap (Chen et al., 2023a) VLFM (Yokoyama et al., 2023) VoxPoser (Huang et al., 2023b) InstructNav (Long et al., 2024) OVL-MAP (Wen et al., 2025) LM-Nav (Shah et al., 2023) RoboHop (Garg et al., 2024) Clio (Maggio et al., 2024) OpenScene (Peng et al., 2023) CLIPFields (Shafiullah et al., 2023) ConceptFusion (Jatavallabhula et al., 2023) CLIP2Scene (Chen et al., 2023c) 3D aware ObjNav (Zhang et al., 2023a)
ConceptGraphs (Gu et al., 2023b) HOV-SG (Werby et al., 2024)